Please help me solve this question :(
The following questions involve either bridge hands (13 cards) or poker hands (5 cards):

a) how many possible poker hands exist?
b) how many poker hands contain all Hearts?
c) how many poker hands contain at least one Ace?
d) how many possible bridge hands contain all spades OR all hearts?
e) how many bridge hands contain exactly 5 Hearts, 5 Diamonds, 2 Spades and one Club?

Answer by Edwin McCravy(20055)

The following questions involve either bridge hands (13 cards) or poker hands (5 cards):
a) how many possible poker hands exist?
C(52,5)

b) how many poker hands contain all Hearts?
C(13,5)

c) how many poker hands contain at least one Ace?
Number of poker hands MINUS the number of no-Ace handa

C(52,5)-C(48,5)

d) how many possible bridge hands contain all spades OR all hearts?
C(13,13)+C(13,13) = 1+1= 2.  (Either all the spades or all the hearts)

e) how many bridge hands contain exactly 5 Hearts, 5 Diamonds, 2 Spades and one Club
C(13,5)*C(13,5)*C(13,2)*C(13,1)

Edwin